OpenStudy (ashleyisakitty):

How does x^1/4 * x^4 equal 17/4? Thanks.

Nnesha (nnesha):

do you mean x^{17}/4 ??

OpenStudy (ashleyisakitty):

Yes :)

Nnesha (nnesha):

\[x^m *x^n =x^{m+n}\] the exponent rule when multiply same bases we should add the exponents :)

Nnesha (nnesha):

so that's how \[\large\rm x^\frac{ 1 }{ 4 } +x^4 = x^{\frac{ 1 }{ 4 }+4}\]

OpenStudy (ashleyisakitty):

Right, I know that. I am confused on how it equates to 17/4.

OpenStudy (bahrom7893):

\[\frac{1}{4}+4 = \frac{1+4*4}{4}\]

Nnesha (nnesha):

find the common denominator 4 is same as 4/1 so \[\frac{ 1 }{ 4 } +\frac{4}{1}\]

OpenStudy (ashleyisakitty):

Thank you!